OpenStudy (anonymous):

simplify and reduce to lowest terms 1/x+4

OpenStudy (anonymous):

\[\frac{ 1 }{ x}+4\]

OpenStudy (loser66):

can you explain me how to get that?

OpenStudy (anonymous):

thats the actuall problem that needs to be simplified and reduced

OpenStudy (loser66):

I don't know how \[\frac{1}{x+4}=\frac{1}{x}+ 4\]

OpenStudy (anonymous):

i thought it would be \[\frac{ 5 }{ x+4 }\]

OpenStudy (jdoe0001):

@sndrod09 do you know how to get an LCD?

OpenStudy (anonymous):

i thinkso

OpenStudy (jdoe0001):

ok, so bearing in mind that \(\large 4=\frac{4}{1}\) then $$ \cfrac{1}{x}+\cfrac{4}{1} = \cfrac{}{\boxed{LCD?}} $$

OpenStudy (anonymous):



OpenStudy (anonymous):

mulitplyby 1?

OpenStudy (anonymous):

lCD =1

OpenStudy (anonymous):

orisit x

OpenStudy (jdoe0001):

I guess I meant the LCM :S

OpenStudy (anonymous):

1x would be LCMright ?

OpenStudy (jdoe0001):

maybe not, it'd be the LCD, anyhow, lemme get the definition correct ehhehe

OpenStudy (jdoe0001):

an yhow, it's lcd, or lowest ... well, a number/term, you DIVIDE by ALL denominators

OpenStudy (jdoe0001):

so, say for the denominators, 3 and 6, a common will be "6", because, 6/3 = 2, and 6/6 =1, now, 5, won't work to produce an integer in this case, so that won't work, 13 is not divisible by 3 or 6, so that won't work either

OpenStudy (anonymous):

right so am i just making up a number for x?

OpenStudy (anonymous):

x can be any real number right ?

OpenStudy (jdoe0001):

well, in this case, yes

OpenStudy (jdoe0001):

http://www.math.com/school/subject1/images/SIU3L3GL.gif they call it the LCM there, but is the same thing, the lcd :|

OpenStudy (anonymous):

withthat being the case the answer shouldbe 5/x

OpenStudy (jdoe0001):

hmm, do you know how to add fractions?

OpenStudy (anonymous):

or\[\frac{ 4x+1 }{ x }\]

OpenStudy (anonymous):

DUETOME MULTIPLYING 4(x)/1(x)

OpenStudy (jdoe0001):

ding :D $$ \cfrac{1}{x}+\cfrac{4}{1} = \cfrac{1+4x}{x} \implies \cfrac{4x+1}{x} $$

OpenStudy (anonymous):

sweetthanks